A company invests equal money into 2 investments. One investment pays 7% interest/yr and the other pays 8% interest/yr. After one year the interest is $3600. How much money was invested into each investment?
-2 investments
-1@ 7%
-1@ 8%
- total interest is $3600

Answer by mananth(16946)

7% .............. $x
8%............... $x
..
Interest = $3,600
7x+8x= 3600 *100
15x=360,000
/15
x=360,000/15
x = $24,000
...
m.ananth@hotmail.ca